# Graphon Quantum Filtering systems

### v1

@@ 5.2 Propagation of chaos
$(\star\star)^{j_{1}}$ $=\mathrm{t}\mathrm{r}\Big{(}\big{[}\frac{1}{cN}\sum_{i_{l}\neq j_{1}}{w}_{ji}A_{ji_{l}}-\frac{1}{c}\sum_{i=1}^{c}{w}_{ji}A_{j}^{\mathbb{E}\big{[}\boldsymbol{\gamma}_{t,i}\big{]}},\boldsymbol{g}_{t,j}\big{]}\boldsymbol{\rho}_{t}^{cN}\Big{)}$
$=\frac{1}{c}\mathrm{t}\mathrm{r}\Big{(}\Big{[}\frac{1}{N}\Big{(}\sum_{i\neq j}\sum_{l=1}^{N}{w}_{ji}A_{ji_{l}}+\sum_{l=2}^{N}{w}_{jj}A_{jj_{l}}\Big{)}-\sum_{i=1}^{c}{w}_{ji}A_{j}^{\mathbb{E}\big{[}\boldsymbol{\gamma}_{t,i}\big{]}},\boldsymbol{g}_{t,j}\Big{]}\boldsymbol{\rho}_{t}^{cN}\Big{)}$
$=\frac{1}{c}\mathrm{t}\mathrm{r}\Big{(}\Big{[}\frac{1}{N}\sum_{i\neq j}\sum_{l=1}^{N}{w}_{ji}A_{ji_{l}}-\sum_{i\neq j}^{c}{w}_{ji}A_{j}^{\mathbb{E}\big{[}\boldsymbol{\gamma}_{t,i}\big{]}}+\frac{1}{N}\sum_{l=2}^{N}{w}_{jj}A_{jj_{l}}-{w}_{jj}A_{j}^{\mathbb{E}\big{[}\boldsymbol{\gamma}_{t,j}\big{]}},\boldsymbol{g}_{t,j}\Big{]}\boldsymbol{\rho}_{t}^{cN}\Big{)}$
$=\frac{1}{c}\mathrm{t}\mathrm{r}\Big{(}\Big{[}\sum_{i\neq j}^{c}{w}_{ji}\Big{(}\frac{1}{N}\sum_{l=1}^{N}A_{ji_{l}}-A_{j}^{\mathbb{E}\big{[}\boldsymbol{\gamma}_{t,i}\big{]}}\Big{)}+{w}_{jj}\Big{(}\frac{1}{N}\sum_{l=2}^{N}A_{jj_{l}}-A_{j}^{\mathbb{E}\big{[}\boldsymbol{\gamma}_{t,j}\big{]}}\Big{)},\boldsymbol{g}_{t,j}\Big{]}\boldsymbol{\rho}_{t}^{cN}\Big{)}.$
@@
    $(\star\star\star)_{1}^{j_{1}}$ $:=\frac{1}{c}\mathrm{t}\mathrm{r}\Big{(}\Big{[}\sum_{i\neq j}^{c}{w}_{ji}\Big{(}\frac{1}{N}\sum_{l=1}^{N}A_{ji_{l}}-A_{j}^{\mathbb{E}\big{[}\boldsymbol{\gamma}_{t,i}\big{]}}\Big{)},\boldsymbol{g}_{t,j}\Big{]}\boldsymbol{\rho}_{t}^{cN}\Big{)}$
@@
    $(\star\star\star)_{2}^{j_{1}}$ $:=\frac{1}{c}\mathrm{t}\mathrm{r}\Big{(}\Big{[}{w}_{jj}\Big{(}\frac{1}{N}\sum_{l=2}^{N}A_{jj_{l}}-A_{j}^{\mathbb{E}\big{[}\boldsymbol{\gamma}_{t,j}\big{]}}\Big{)},\boldsymbol{g}_{t,j}\Big{]}\boldsymbol{\rho}_{t}^{cN}\Big{)}.$
We focus below on the inter-block term $(\star\star\star)_{1}^{j_{1}}$. The treatment of the intra-block term $(\star\star\star)_{2}^{j_{1}}$ follows from the same arguments used for the complete graph case in [43].
Estimate $(\star\star\star)_{1}^{j_{1}}$
@@
$\delta^{j_{l^{\prime}},N}_{t}$ $:=\frac{1}{N}\sum_{l^{\prime\prime}=1}^{N}\gamma_{t,j_{l^{\prime\prime}}}-\mathbb{E}[\gamma_{t,j_{l^{\prime}}}],\quad A_{j}^{\mathbb{E}[\gamma_{t,j_{l^{\prime}}}]}:=\frac{1}{N}\sum_{l^{\prime\prime}=1}^{N}A_{j}^{\gamma_{t,j_{l^{\prime\prime}}}}-A_{j}^{\delta_{t}^{j_{l^{\prime}}},N}.$
Since $\mathbb{E}\big{[}\delta_{t}^{j,l,N}\big{]}=0$, and by law of large numbers, we obtain
$\mathbb{E}\Big{[}\|\delta_{t}^{j,l,N}\|_{2}^{2}\Big{]}$ $\stackrel{{\scriptstyle}}{{=}}\mathbb{V}\Big{[}\|\delta_{t}^{j,l,N}\|_{2}^{2}\Big{]}\leq\frac{1}{N}\mathbb{V}\Big{[}\|\gamma_{t,j_{l}}\|_{2}^{2}\Big{]}.$ (5.4)
Then,
@@
Estimate $(\star\star\star)_{11}^{j_{1}}$
@@
Estimate $(\star\star\star)_{12}^{j_{1}}$
@@
$(\alpha_{2})^{i_{l}}$ $=\langle\Psi_{t}^{cN},\boldsymbol{g}_{t,j}\boldsymbol{g}_{t,i_{l}}\big{(}A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big{)}\boldsymbol{g}_{t,j}\boldsymbol{\gamma}_{t,i_{l}}\Psi_{t}^{cN}\rangle,$ $\quad(\beta_{2})^{i_{l}}$ $=\langle\Psi_{t}^{cN},\boldsymbol{g}_{t,j}\boldsymbol{\gamma}_{t,i_{l}}\big{(}A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big{)}\boldsymbol{\gamma}_{t,j}\boldsymbol{g}_{t,i_{l}}\Psi_{t}^{cN}\rangle,$
$(\alpha_{3})^{i_{l}}$ $=\langle\Psi_{t}^{cN},\boldsymbol{g}_{t,j}\boldsymbol{\gamma}_{t,i_{l}}\big{(}A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big{)}\boldsymbol{g}_{t,j}\boldsymbol{g}_{t,i_{l}}\Psi_{t}^{cN}\rangle,$ $\quad(\beta_{3})^{i_{l}}$ $=\langle\Psi_{t}^{cN},\boldsymbol{g}_{t,j}\boldsymbol{\gamma}_{t,i_{l}}\big{(}A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big{)}\boldsymbol{g}_{t,j}\boldsymbol{\gamma}_{t,i_{l}}\Psi_{t}^{cN}\rangle,$
$(\alpha_{4})^{i_{l}}$ $=\langle\Psi_{t}^{cN},\boldsymbol{g}_{t,j}\boldsymbol{\gamma}_{t,i_{l}}\big{(}A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big{)}\boldsymbol{g}_{t,j}\boldsymbol{\gamma}_{t,i_{l}}\Psi_{t}^{cN}\rangle,$ $\quad(\beta_{4})^{i_{l}}$ $=\langle\Psi_{t}^{cN},\boldsymbol{g}_{t,j}\boldsymbol{\gamma}_{t,i_{l}}\big{(}A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big{)}\boldsymbol{g}_{t,j}\boldsymbol{g}_{t,i_{l}}\Psi_{t}^{cN}\rangle,$
$(\alpha_{5})^{i_{l}}$ $=\langle\Psi_{t}^{cN},\boldsymbol{\gamma}_{t,j}\boldsymbol{g}_{t,i_{l}}\big{(}A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big{)}\boldsymbol{g}_{t,j}\boldsymbol{g}_{t,i_{l}}\Psi_{t}^{cN}\rangle,$ $\quad(\beta_{5})^{i_{l}}$ $=\langle\Psi_{t}^{cN},\boldsymbol{g}_{t,j}\boldsymbol{g}_{t,i_{l}}\big{(}A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big{)}\boldsymbol{\gamma}_{t,j}\boldsymbol{\gamma}_{t,i_{l}}\Psi_{t}^{cN}\rangle,$
$(\alpha_{6})^{i_{l}}$ $=\langle\Psi_{t}^{cN},\boldsymbol{\gamma}_{t,j}\boldsymbol{g}_{t,i_{l}}\big{(}A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big{)}\boldsymbol{g}_{t,j}\boldsymbol{\gamma}_{t,i_{l}}\Psi_{t}^{cN}\rangle,$ $\quad(\beta_{6})^{i_{l}}$ $=\langle\Psi_{t}^{cN},\boldsymbol{g}_{t,j}\boldsymbol{g}_{t,i_{l}}\big{(}A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big{)}\boldsymbol{\gamma}_{t,j}\boldsymbol{g}_{t,i_{l}}\Psi_{t}^{cN}\rangle,$
$(\alpha_{7})^{i_{l}}$ $=\langle\Psi_{t}^{cN},\boldsymbol{\gamma}_{t,j}\boldsymbol{\gamma}_{t,i_{l}}\big{(}A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big{)}\boldsymbol{g}_{t,j}\boldsymbol{g}_{t,i_{l}}\Psi_{t}^{cN}\rangle,$ $\quad(\beta_{7})^{i_{l}}$ $=\langle\Psi_{t}^{cN},\boldsymbol{g}_{t,j}\boldsymbol{g}_{t,i_{l}}\big{(}A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big{)}\boldsymbol{g}_{t,j}\boldsymbol{\gamma}_{t,i_{l}}\Psi_{t}^{cN}\rangle,$
$(\alpha_{8})^{i_{l}}$ $=\langle\Psi_{t}^{cN},\boldsymbol{\gamma}_{t,j}\boldsymbol{\gamma}_{t,i_{l}}\big{(}A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big{)}\boldsymbol{g}_{t,j}\boldsymbol{\gamma}_{t,i_{l}}\Psi_{t}^{cN}\rangle,$ $\quad(\beta_{8})^{i_{l}}$ $=\langle\Psi_{t}^{cN},\boldsymbol{g}_{t,j}\boldsymbol{g}_{t,i_{l}}\big{(}A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big{)}\boldsymbol{g}_{t,j}\boldsymbol{g}_{t,i_{l}}\Psi_{t}^{cN}\rangle.$
@@
Applying the technical Lemma 3,
@@
$\mathbb{E}\Big{[}(\triangle_{2}^{j,i_{l}})\Big{]}$ $\stackrel{{\scriptstyle\text{C-S}}}{{\leq}}\sqrt{\mathbb{E}\Big{[}\Big{\|}A_{ji_{l}}\hat{\tau_{1}h_{i}}\hat{\tau_{1}h_{j}}\boldsymbol{\gamma}_{t,j}\boldsymbol{\gamma}_{t,i_{l}}\Psi_{t}^{cN}\Big{\|}_{2}^{2}\Big{]}}\sqrt{\mathbb{E}\Big{[}\Big{\|}\hat{h}^{-1}_{i}\hat{h}^{-1}_{j}\boldsymbol{g}_{t,j}\boldsymbol{g}_{t,i_{l}}\Psi_{t}^{cN}\Big{\|}_{2}^{2}\Big{]}}.$
@@
$\mathbb{E}\Big{[}\Big{\|}\hat{h}^{-1}_{i}\hat{h}^{-1}_{j}\boldsymbol{g}_{t,j}\boldsymbol{g}_{t,i_{l}}\Psi_{t}^{cN}\Big{\|}_{2}^{2}\Big{]}=\mathbb{E}\Big{[}\langle\Psi_{t}^{cN},\hat{h}_{i}^{-2}\hat{h}^{-2}_{j}\boldsymbol{g}_{t,i}\boldsymbol{g}_{t,i_{l}}\Psi^{cN}\rangle\Big{]}$
$\quad\stackrel{{\scriptstyle\text{C-S}}}{{\leq}}\mathbb{E}\Big{[}\|\hat{f}_{i}^{-1}\hat{f}_{j}^{-1}\|_{\text{op}}\Big{\|}\Psi_{t}^{cN}\boldsymbol{g}_{t,i}\Big{\|}_{2}^{2}\Big{\|}\Psi_{t}^{cN}\boldsymbol{g}_{t,j_{l}}\Big{\|}_{2}^{2}\Big{]}\leq C\mathbb{E}\Big{[}\mathcal{D}_{N}^{i}(t)\mathcal{D}_{N}^{j}(t)\Big{]}\leq C\underbrace{\mathbb{E}\Big{[}\mathcal{D}_{N}(t)\Big{]}}_{{x}},$

### v3

@@ 5.2 Propagation of chaos
$(\star\star)^{j_{1}}$ $=\mathrm{t}\mathrm{r}\Big(\big[\frac{1}{cN}\sum_{i_{l}\neq j_{1}}{w}_{ji}A_{ji_{l}}-\frac{1}{c}\sum_{i=1}^{c}{w}_{ji}A_{j}^{\mathbb{E}_{\mathbb{P}}\big[\boldsymbol{\gamma}_{t,i}\big]},\boldsymbol{g}_{t,j}\big]\boldsymbol{\rho}_{t}^{cN}\Big)$
$=\frac{1}{c}\mathrm{t}\mathrm{r}\Big(\Big[\frac{1}{N}\Big(\sum_{i\neq j}\sum_{l=1}^{N}{w}_{ji}A_{ji_{l}}+\sum_{l=2}^{N}{w}_{jj}A_{jj_{l}}\Big)-\sum_{i=1}^{c}{w}_{ji}A_{j}^{\mathbb{E}_{\mathbb{P}}\big[\boldsymbol{\gamma}_{t,i}\big]},\boldsymbol{g}_{t,j}\Big]\boldsymbol{\rho}_{t}^{cN}\Big)$
$=\frac{1}{c}\!\mathrm{t}\mathrm{r}\Big(\Big[\frac{1}{N}\sum_{i\neq j}\sum_{l=1}^{N}{w}_{ji}A_{ji_{l}}-\sum_{i\neq j}^{c}{w}_{ji}A_{j}^{\mathbb{E}_{\mathbb{P}}\big[\boldsymbol{\gamma}_{t,i}\big]}\!+\!\frac{1}{N}\sum_{l=2}^{N}{w}_{jj}A_{jj_{l}}-{w}_{jj}A_{j}^{\mathbb{E}_{\mathbb{P}}\big[\boldsymbol{\gamma}_{t,j}\big]},\boldsymbol{g}_{t,j}\Big]\boldsymbol{\rho}_{t}^{cN}\Big)$
$=\frac{1}{c}\mathrm{t}\mathrm{r}\Big(\Big[\sum_{i\neq j}^{c}{w}_{ji}\Big(\frac{1}{N}\sum_{l=1}^{N}A_{ji_{l}}-A_{j}^{\mathbb{E}\big[\boldsymbol{\gamma}_{t,i}\big]}\Big)+{w}_{jj}\Big(\frac{1}{N}\sum_{l=2}^{N}A_{jj_{l}}-A_{j}^{\mathbb{E}_{\mathbb{P}}\big[\boldsymbol{\gamma}_{t,j}\big]}\Big),\boldsymbol{g}_{t,j}\Big]\boldsymbol{\rho}_{t}^{cN}\Big).$
@@
    $(\star\star\star)_{1}^{j_{1}}$ $:=\frac{1}{c}\mathrm{t}\mathrm{r}\Big(\Big[\sum_{i\neq j}^{c}{w}_{ji}\Big(\frac{1}{N}\sum_{l=1}^{N}A_{ji_{l}}-A_{j}^{\mathbb{E}_{\mathbb{P}}\big[\boldsymbol{\gamma}_{t,i}\big]}\Big),\boldsymbol{g}_{t,j}\Big]\boldsymbol{\rho}_{t}^{cN}\Big)$
@@
    $(\star\star\star)_{2}^{j_{1}}$ $:=\frac{1}{c}\mathrm{t}\mathrm{r}\Big(\Big[{w}_{jj}\Big(\frac{1}{N}\sum_{l=2}^{N}A_{jj_{l}}-A_{j}^{\mathbb{E}_{\mathbb{P}}\big[\boldsymbol{\gamma}_{t,j}\big]}\Big),\boldsymbol{g}_{t,j}\Big]\boldsymbol{\rho}_{t}^{cN}\Big).$
We focus below on the inter-block term $(\star\star\star)_{1}^{j_{1}}$. The treatment of the intra-block term $(\star\star\star)_{2}^{j_{1}}$ follows the same arguments used for the homogeneous case in [47].
Estimate $(\star\star\star)_{1}^{j_{1}}$.
@@
$\delta^{j_{l^{\prime}},N}_{t}$ $:=\frac{1}{N}\sum_{l^{\prime\prime}=1}^{N}\gamma_{t,j_{l^{\prime\prime}}}-\mathbb{E}_{\mathbb{P}}[\gamma_{t,j_{l^{\prime}}}],\quad A_{j}^{\mathbb{E}_{\mathbb{P}}[\gamma_{t,j_{l^{\prime}}}]}:=\frac{1}{N}\sum_{l^{\prime\prime}=1}^{N}A_{j}^{\gamma_{t,j_{l^{\prime\prime}}}}-A_{j}^{\delta_{t}^{j_{l^{\prime}}},N}.$
Since $\mathbb{E}_{\mathbb{P}}\big[\delta_{t}^{j,l,N}\big]=0$, and by law of large numbers, we obtain
$\mathbb{E}_{\mathbb{P}}\Big[\|\delta_{t}^{j,l,N}\|_{2}^{2}\Big]$ $\stackrel{{\scriptstyle}}{{=}}\mathbb{V}\Big[\|\delta_{t}^{j,l,N}\|_{2}^{2}\Big]\leq\frac{1}{N}\mathbb{V}\Big[\|\gamma_{t,j_{l}}\|_{2}^{2}\Big],$ (14)
where $\mathbb{V}$ is the variance under $\mathbb{P}$, then,
@@
Estimate $(\star\star\star)_{11}^{j_{1}}$.
@@
Estimate $(\star\star\star)_{12}^{j_{1}}$.
@@
$(\alpha_{2})^{i_{l}}=\langle\Psi_{t}^{cN},\boldsymbol{g}_{t,j}\boldsymbol{g}_{t,i_{l}}\big(A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big)\boldsymbol{g}_{t,j}\boldsymbol{\gamma}_{t,i_{l}}\Psi_{t}^{cN}\rangle,$ $\ (\beta_{2})^{i_{l}}=\langle\Psi_{t}^{cN},\boldsymbol{g}_{t,j}\boldsymbol{\gamma}_{t,i_{l}}\big(A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big)\boldsymbol{\gamma}_{t,j}\boldsymbol{g}_{t,i_{l}}\Psi_{t}^{cN}\rangle,$
$(\alpha_{3})^{i_{l}}=\langle\Psi_{t}^{cN},\boldsymbol{g}_{t,j}\boldsymbol{\gamma}_{t,i_{l}}\big(A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big)\boldsymbol{g}_{t,j}\boldsymbol{g}_{t,i_{l}}\Psi_{t}^{cN}\rangle,$ $\ (\beta_{3})^{i_{l}}=\langle\Psi_{t}^{cN},\boldsymbol{g}_{t,j}\boldsymbol{\gamma}_{t,i_{l}}\big(A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big)\boldsymbol{g}_{t,j}\boldsymbol{\gamma}_{t,i_{l}}\Psi_{t}^{cN}\rangle,$
$(\alpha_{4})^{i_{l}}=\langle\Psi_{t}^{cN},\boldsymbol{g}_{t,j}\boldsymbol{\gamma}_{t,i_{l}}\big(A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big)\boldsymbol{g}_{t,j}\boldsymbol{\gamma}_{t,i_{l}}\Psi_{t}^{cN}\rangle,$ $\ (\beta_{4})^{i_{l}}=\langle\Psi_{t}^{cN},\boldsymbol{g}_{t,j}\boldsymbol{\gamma}_{t,i_{l}}\big(A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big)\boldsymbol{g}_{t,j}\boldsymbol{g}_{t,i_{l}}\Psi_{t}^{cN}\rangle,$
$(\alpha_{5})^{i_{l}}=\langle\Psi_{t}^{cN},\boldsymbol{\gamma}_{t,j}\boldsymbol{g}_{t,i_{l}}\big(A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big)\boldsymbol{g}_{t,j}\boldsymbol{g}_{t,i_{l}}\Psi_{t}^{cN}\rangle,$ $\ (\beta_{5})^{i_{l}}=\langle\Psi_{t}^{cN},\boldsymbol{g}_{t,j}\boldsymbol{g}_{t,i_{l}}\big(A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big)\boldsymbol{\gamma}_{t,j}\boldsymbol{\gamma}_{t,i_{l}}\Psi_{t}^{cN}\rangle,$
$(\alpha_{6})^{i_{l}}=\langle\Psi_{t}^{cN},\boldsymbol{\gamma}_{t,j}\boldsymbol{g}_{t,i_{l}}\big(A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big)\boldsymbol{g}_{t,j}\boldsymbol{\gamma}_{t,i_{l}}\Psi_{t}^{cN}\rangle,$ $\ (\beta_{6})^{i_{l}}=\langle\Psi_{t}^{cN},\boldsymbol{g}_{t,j}\boldsymbol{g}_{t,i_{l}}\big(A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big)\boldsymbol{\gamma}_{t,j}\boldsymbol{g}_{t,i_{l}}\Psi_{t}^{cN}\rangle,$
$(\alpha_{7})^{i_{l}}=\langle\Psi_{t}^{cN},\boldsymbol{\gamma}_{t,j}\boldsymbol{\gamma}_{t,i_{l}}\big(A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big)\boldsymbol{g}_{t,j}\boldsymbol{g}_{t,i_{l}}\Psi_{t}^{cN}\rangle,$ $\ (\beta_{7})^{i_{l}}=\langle\Psi_{t}^{cN},\boldsymbol{g}_{t,j}\boldsymbol{g}_{t,i_{l}}\big(A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big)\boldsymbol{g}_{t,j}\boldsymbol{\gamma}_{t,i_{l}}\Psi_{t}^{cN}\rangle,$
$(\alpha_{8})^{i_{l}}=\langle\Psi_{t}^{cN},\boldsymbol{\gamma}_{t,j}\boldsymbol{\gamma}_{t,i_{l}}\big(A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big)\boldsymbol{g}_{t,j}\boldsymbol{\gamma}_{t,i_{l}}\Psi_{t}^{cN}\rangle,$ $\ (\beta_{8})^{i_{l}}=\langle\Psi_{t}^{cN},\boldsymbol{g}_{t,j}\boldsymbol{g}_{t,i_{l}}\big(A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big)\boldsymbol{g}_{t,j}\boldsymbol{g}_{t,i_{l}}\Psi_{t}^{cN}\rangle.$
@@
Applying the technical Lemma 5.3,
@@
$\mathbb{E}_{\mathbb{P}}\Big[(\triangle_{2}^{j,i_{l}})\Big]$ $\stackrel{{\scriptstyle\text{C-S}}}{{\leq}}\sqrt{\mathbb{E}_{\mathbb{P}}\Big[\Big\|A_{ji_{l}}\hat{\tau_{1}h_{i}}\hat{\tau_{1}h_{j}}\boldsymbol{\gamma}_{t,j}\boldsymbol{\gamma}_{t,i_{l}}\Psi_{t}^{cN}\Big\|_{2}^{2}\Big]}\sqrt{\mathbb{E}_{\mathbb{P}}\Big[\Big\|\hat{h}^{-1}_{i}\hat{h}^{-1}_{j}\boldsymbol{g}_{t,j}\boldsymbol{g}_{t,i_{l}}\Psi_{t}^{cN}\Big\|_{2}^{2}\Big]}.$
@@
$\mathbb{E}_{\mathbb{P}}\Big[\Big\|\hat{h}^{-1}_{i}\hat{h}^{-1}_{j}\boldsymbol{g}_{t,j}\boldsymbol{g}_{t,i_{l}}\Psi_{t}^{cN}\Big\|_{2}^{2}\Big]=\mathbb{E}_{\mathbb{P}}\Big[\langle\Psi_{t}^{cN},\hat{h}_{i}^{-2}\hat{h}^{-2}_{j}\boldsymbol{g}_{t,i}\boldsymbol{g}_{t,i_{l}}\Psi^{cN}\rangle\Big]$
$\quad\stackrel{{\scriptstyle\text{C-S}}}{{\leq}}\mathbb{E}_{\mathbb{P}}\Big[\|\hat{f}_{i}^{-1}\hat{f}_{j}^{-1}\|_{\text{op}}\Big\|\Psi_{t}^{cN}\boldsymbol{g}_{t,i}\Big\|_{2}^{2}\Big\|\Psi_{t}^{cN}\boldsymbol{g}_{t,j_{l}}\Big\|_{2}^{2}\Big]\leq C\,\mathbb{E}_{\mathbb{P}}\Big[\mathcal{D}_{N}^{i}(t)\mathcal{D}_{N}^{j}(t)\Big]\leq C\underbrace{\mathbb{E}_{\mathbb{P}}\Big[\mathcal{D}_{N}(t)\Big]}_{{x}},$
$\mathbb{E}_{\mathbb{P}}\Big[\Big\|A_{ji_{l}}\hat{\tau_{1}h_{i}}\hat{\tau_{1}h_{j}}\boldsymbol{\gamma}_{t,j}\boldsymbol{\gamma}_{t,i_{l}}\Psi_{t}^{cN}\Big\|_{2}^{2}\Big]$ $\leq\!\mathbb{E}_{\mathbb{P}}\!\Big[\langle\Psi_{t}^{cN},\hat{\tau_{1}h_{i}}\hat{\tau_{1}h_{j}}\boldsymbol{\gamma}_{t,i_{l}}\boldsymbol{\gamma}_{t,j}A_{ji_{l}}^{2}\hat{\tau_{1}h_{i}}\hat{\tau_{1}h_{j}}\boldsymbol{\gamma}_{t,i_{l}}\boldsymbol{\gamma}_{t,j}\Psi_{t}^{cN}\rangle\Big]$
$\leq\|a\|_{2}\,\mathbb{E}_{\mathbb{P}}\Big[\langle\Psi_{t}^{cN},\hat{\tau_{1}h_{i}^{2}}\Psi_{t}^{cN}\rangle\langle\Psi_{t}^{cN},\hat{\tau_{1}h_{j}^{2}}\Psi_{t}^{cN}\rangle\Big]$
$\leq\|a\|_{2}\,\mathbb{E}_{\mathbb{P}}\Big[\Big(\langle\Psi_{t}^{cN},\hat{f_{i}}\Psi_{t}^{cN}\rangle+\frac{1}{N}\Big)\Big(\langle\Psi_{t}^{cN},\hat{f_{j}}\Psi_{t}^{cN}\rangle+\frac{1}{N}\Big)\Big]$
$\leq\|a\|_{2}\,\mathbb{E}_{\mathbb{P}}\Big[\Big(\mathcal{D}^{i}_{N}(t)+\frac{1}{N}\Big)\Big(\mathcal{D}^{j}_{N}(t)+\frac{1}{N}\Big)\Big]$
$\leq\|a\|_{2}\,\mathbb{E}_{\mathbb{P}}\Big[\Big(\mathcal{D}_{N}(t)+\frac{1}{N}\Big)^{2}\Big]\leq\|a\|_{2}\underbrace{\mathbb{E}_{\mathbb{P}}\Big[\Big(\mathcal{D}_{N}(t)+\frac{3}{N}\Big)\Big]}_{{y}}.$
Hence,
$\mathbb{E}_{\mathbb{P}}\Big[(\triangle_{2}^{j,i_{l}})\Big]$ $\stackrel{{\scriptstyle\sqrt{xy}\leq\frac{x+y}{2}}}{{\leq}}C\Big(\mathbb{E}_{\mathbb{P}}\Big[\mathcal{D}_{N}(t)\Big]+\frac{3}{2N}\Big).$
Estimate $(\triangle_{3}^{j,i_{l}})$.
We have
$\mathbb{E}_{\mathbb{P}}\big[{(\triangle_{3}^{j,i_{l}})}\big]$ $=\mathbb{E}_{\mathbb{P}}\Big[\Big|\langle\Psi_{t}^{cN},\boldsymbol{\gamma}_{t,j}\boldsymbol{g}_{t,i_{l}}\big(A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big)\boldsymbol{g}_{t,j}\boldsymbol{g}_{t,i_{l}}\Psi_{t}^{cN}\rangle\Big|\Big]$
$\stackrel{{\scriptstyle\text{C-S}}}{{\leq}}\mathbb{E}_{\mathbb{P}}\Big[\Big\|\boldsymbol{g}_{t,j}\boldsymbol{g}_{t,i_{l}}\Psi_{t}^{cN}\Big\|_{2}\Big\|\big(A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big)\boldsymbol{\gamma}_{t,j}\boldsymbol{g}_{t,i_{l}}\Psi_{t}^{cN}\Big\|_{2}\Big]$
$\leq\mathbb{E}_{\mathbb{P}}\Big[\Big\|\boldsymbol{g}_{t,j}\boldsymbol{g}_{t,i_{l}}\Psi_{t}^{cN}\Big\|_{2}\Big\|\boldsymbol{\gamma}_{t,j}\big(A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big)^{2}\boldsymbol{\gamma}_{t,j}\Psi_{t}^{cN}\Big\|_{\text{op}}\Big\|\boldsymbol{g}_{t,i_{l}}\Psi_{t}^{cN}\Big\|_{2}\Big]$
$\leq C\,\mathbb{E}_{\mathbb{P}}\Big[\mathcal{D}_{N}^{i_{l}}(t)+\mathcal{D}_{N}^{j_{1}}(t)\Big]\leq C\,\mathbb{E}_{\mathbb{P}}\Big[\mathcal{D}_{N}(t)\Big],$
where we used the bound $\Big\|\boldsymbol{\gamma}_{t,j}\big(A_{ji_{l}}-A^{\gamma_{t,i_{l}}}_{j}\big)^{2}\boldsymbol{\gamma}_{t,j}\Psi_{t}^{cN}\Big\|_{\text{op}}\leq C.$
We thus have
$\mathbb{E}_{\mathbb{P}}\Big[\Big|(\star\star\star)_{12}^{j_{1}}\Big|\Big]$ $\leq C\Big(\mathbb{E}_{\mathbb{P}}[\mathcal{D}_{N}(t)]+\frac{1}{N}\Big),$